OpenStudy (anonymous):

Find the quotient. Simplify your answer. 2b/9 (divided) -5c/3

OpenStudy (lalaly):

\[\frac{2b}{9} \div \frac{-5c}{3}=\frac{2b}{9} \times \frac{3}{-5c}\]

OpenStudy (anonymous):

Is that the answer?

OpenStudy (lalaly):

no u simplify it

OpenStudy (anonymous):

I don't know how to. That's my problem.

OpenStudy (anonymous):

6b/35c